A graduate student in chemistry needs 172 liters of a 41% saline solution. How many liters of an 18% saline solution must be mixed with a 61% saline solutionto obtain the required new solution?
Answer by stanbon(75887)

A graduate student in chemistry needs 172 liters of a 41% saline solution. How many liters of an 18% saline solution must be mixed with a 61% saline solutionto obtain the required new solution?
---------------------------
Equation:
salt + salt = salt
0.61x + 0.18(172-x) = 0.41(172)
61x + 18*172 - 18x = 41*172
43x = 24*172
x = 96 liters (amt. of 61% solution in the mix)
172- x = 76 liters ( amt. of 18% solution in the mix)
========================================================
Cheers,
Stan H.